OpenStudy (anonymous):

8x+4=4(2x+1) is this infinite real solutions or no real solutions or one real solution

OpenStudy (jennyrlz):

What math is this?

OpenStudy (jennyrlz):

I am not really confident but i will give it a go

OpenStudy (anonymous):

okay

OpenStudy (jennyrlz):

8x+4=4(2x+1) distribute 8x+4=8x+4 based on this i would think there is only one real solution

OpenStudy (jennyrlz):

but i dont even know what math is this or what you are expected to do so sorry :/

OpenStudy (jennyrlz):

this is*

OpenStudy (anonymous):

yea i think thts it i mean this is pre algebra

OpenStudy (jennyrlz):

oh well then i think i am correct xD, but dont take my word for it :P

OpenStudy (anonymous):

oh pls i wil

OpenStudy (jennyrlz):

word of advice dont take pre-calc, i am in it and.... it is not fun ; ;

OpenStudy (anonymous):

oh

OpenStudy (jennyrlz):

but if you think you can handle it dont be discouraged xD

OpenStudy (anonymous):

idk

OpenStudy (jennyrlz):

You need help with anything else?

OpenStudy (anonymous):

i do

OpenStudy (jennyrlz):

shoot i will try my best :P

OpenStudy (anonymous):

3x-10=4x-13 infinte real solution or no real solution or on real solution

OpenStudy (jennyrlz):

x=3

OpenStudy (jennyrlz):

i just simplified

OpenStudy (jennyrlz):

-10+13

OpenStudy (jennyrlz):

4x-3x

OpenStudy (jennyrlz):

one real solution...

OpenStudy (jennyrlz):

makes me think, does the first question you asked have no solution because when you simplify it, it is: 0=0

OpenStudy (jennyrlz):

8x-8x=0 4-4=0 0=0

OpenStudy (anonymous):

yea

OpenStudy (jennyrlz):

i hope i am helping and not confusing you xD

OpenStudy (anonymous):

yea i guess u r

OpenStudy (jennyrlz):

sec

OpenStudy (anonymous):

okay

OpenStudy (jennyrlz):

ok i am back :)

OpenStudy (anonymous):

okay

OpenStudy (jennyrlz):

well anymore questions?

OpenStudy (anonymous):

yea hold on

OpenStudy (anonymous):

11-2x+4=8+5x is this one solution or infinte solution or no solution

OpenStudy (anonymous):

@jennyrlz

OpenStudy (jennyrlz):

ok i am here xD

OpenStudy (anonymous):

okay then find the answer pls

OpenStudy (anonymous):

i found one solution

OpenStudy (jennyrlz):

hmm simplifying we get 7=7x simplify that 1=x

OpenStudy (jennyrlz):

what did you find?

OpenStudy (anonymous):

yea i got 1 too

OpenStudy (jennyrlz):

yay xD

OpenStudy (anonymous):

thnx

OpenStudy (jennyrlz):

is that all?

OpenStudy (anonymous):

yea i think so

OpenStudy (jennyrlz):

dont forget to medal xP

OpenStudy (anonymous):

yea srry i forgot

OpenStudy (jennyrlz):

have good day :P, if i am ever on @ me and ill see what i can do xD

OpenStudy (anonymous):

then lets message each other

OpenStudy (anonymous):

@jennyrlz